OpenStudy (anonymous):

-4(x + 10) - 6 = -3(x - 2). I need help solving this please.

OpenStudy (anonymous):

do cross multiply

OpenStudy (anonymous):



OpenStudy (anonymous):

Once you cross multiply, you can combine like terms.

OpenStudy (anonymous):

so then it would be -40 times -6 = 2 times -6? or something like that

OpenStudy (anonymous):

sorry im having alot of trouble focusing cri

OpenStudy (anonymous):

\[-4(x+10)-6=-3(x-2)\]\[-4x-40-6=-3x+6\]

OpenStudy (anonymous):

add like terma as in the x variable

OpenStudy (anonymous):

terms**

OpenStudy (anonymous):

\[-4x-46=-3x+6\]

OpenStudy (anonymous):

Now try to get the like terms grouped together.

OpenStudy (anonymous):

Are you getting it so far ? @misty_fox

OpenStudy (anonymous):

Yes im just trying writing it down on paper, one sec

OpenStudy (anonymous):

-46?

OpenStudy (anonymous):

Ya. -40-6 = -46. I combined those numbers in the second step.

OpenStudy (anonymous):

So what's next ? @misty_fox

OpenStudy (anonymous):

would that be the final answer tho? cuz that is one of the options

OpenStudy (anonymous):

You would have to bring the x terms on one side of the equation after that 3rd step. Do you know how to do that?

OpenStudy (anonymous):

Yea i did that and i got -47 but its closest to -46 and i might have did it wrong but idk

OpenStudy (anonymous):

\[-4x-46=-3x+6\]\[-4x+3x=46+6\]\[-x=52\]\[x=-52\]

OpenStudy (anonymous):

Is -52 one of the answer choices?

OpenStudy (anonymous):

I just got that too ^^^

OpenStudy (anonymous):

Okay great! :)

OpenStudy (anonymous):

Did you get that @misty_fox?

OpenStudy (anonymous):

Yes sorry i was afk, thank you very much! :)

OpenStudy (anonymous):

You're welcome!